OpenStudy (anonymous):

Find the roots of the polynomial equation. -x3 + 5x2 - 11x + 55 = 0

OpenStudy (anonymous):

\[ -x^3+5 x^2-11 x+55=(5-x) \left(x^2+11\right)=0 \] x=5 is the only real root

OpenStudy (anonymous):

Thank you